Assignment 4-4 Week 4 Problems G Help Save & Exit Submit Saved Check my work Problem 5-20 CVP Applications: Break-Even Analysis; Cost Structure; Target Sales [LO5-1, LO5-3, LO5- 4, LO5-5, LO5-6, LO5-8] ints Northwood Company manufactures basketballs. The company has a ball that sells for $38. At present, the ball is manufactured in a small plant that relies heavily on direct labor workers. Thus, variable expenses are high, totaling $28.00 per ball, of which 74% is direct labor cost. Skipped Last year, the company sold 30,000 of these balls, with the following results: Print Sales (30,000 balls) Variable expenses Contribution margin Fixed expenses Net operating income s 1,140,900 eferences 300,000 210,800 $ 90,900 Required 1. Compute (a) last years CM ratio and the break-even point in balls, and (b) the degree of operating leverage at last years sales level. 2. Due to an increase in labor rates, the company estimates that next years variable expenses will increase by $3.00 per ball. If this change takes place and the selling price per ball remains constant at $38.00, what will be next years CM ratio and the break-even point in balls? 3. Refer to the data in (2) above. If the expected change in variable expenses takes place, how many balls will have to be sold next Prey 1 of 2 Next >

media%2F715%2F715779a9-55c5-4b73-b1aa-91

media%2F7f9%2F7f937690-ac77-49ab-97ca-c9

media%2Fe2c%2Fe2c98e29-243b-4ee2-b9ab-65

media%2F51b%2F51be3dcc-2e2c-4eef-ae75-2e

media%2F331%2F33107711-b29e-465d-8620-01

media%2F76d%2F76d4f71a-e270-4f39-aa33-ab

media%2F29b%2F29b27cf5-bd91-46aa-bc82-e6

media%2F682%2F682607c2-7fdd-4abb-b594-63
Answer #1
1
a. Contribution margin ratio=contribution margin per unit/Sales price per unit=10/38=0.26
Contribution margin per unit=Sales price per ball-Variable cost per ball=38-28=10
Break-even point in balls=Fixed expenses/Contribution margin per unit
Break-even point in balls=210000/10=21000 balls
b. Degree of operating leverage=Contribution margin/Operating income=300000/90000=3.33
2 Contribution margin ratio=contribution margin per unit/Sales price per unit
Contribution margin per unit=Sales price per ball-Variable cost per ball=38-(28+3)=7
Contribution margin ratio=7/38=0.18
Break-even point in balls=Fixed expenses/Contribution margin per unit=210000/7=30000 balls
3 Balls required to attain desired profit=(Desired profit+Fixed expenses)/Contribution margin per unit=(90000+210000)/7=42857 balls
4 Assume selling price per ball=x
Contribution margin ratio=contribution margin per unit/Sales price per unit
0.26=(x-31)/x
0.26x=x-31
x-0.26x=31
0.74x=31
x=$41.89
5 Contribution margin ratio=contribution margin per unit/Sales price per unit
Contribution margin per unit=38-[28*(1-0.2632)]=38-20.63=17.37
Contribution margin ratio=contribution margin per unit/Sales price per unit=17.37/38=0.46
Break-even point in balls=Fixed expenses/Contribution margin per unit=(210000*2)/17.37=24179.62 balls
6
a. Balls required to attain desired profit=(Desired profit+Fixed expenses)/Contribution margin per unit=(90000+420000)/17.37=29360.97 balls
b. Contribution income statement
Sales (30000*38) 1140000
Less: Variable expenses
(30000*20.63) 618900
Contribution margin 521100
Less: Fixed expenses 420000
Net operating income 101100
Degree of operating leverage=Contribution margin/Operating income=521100/101100=5.15